OpenStudy (nikki_skellington):

How do banks affect individuals, businesses, government, and society as a whole?

OpenStudy (will.h):

that is my specialist :)

OpenStudy (will.h):

do you need a paragraph or do you need examples?

OpenStudy (nikki_skellington):

a paragraph plz :D

OpenStudy (nikki_skellington):

oops i mean examples

OpenStudy (will.h):

i will add both

OpenStudy (nikki_skellington):

ok thx

OpenStudy (will.h):

Banks are the engine of each country's economy, because Banks hold nearly all of the investment and Banks control the stock market and currencies. So in many ways we may find that banks are what keeps your investment save and in many times when you deposit or withdrawal money you are the winner. here are some of many services that banks offer: 1- saving accounts. Lets say that you have 50,000$ and one of many options allowed to you is depositing that money in a specific bank, in many ways banks have different interest rate but lets say that the interest rate offered by a specific bank is 15% annually so after 5 years how much will your money be? Using Exponential function we can determine that. f(x) = p(1+r)^x where p = the initial deposit r= interest rate x= time Therefore f(x) = 50,000(1+15/100)^5 Therefore after 5 years the deposit would turn to = 100567.859$ Therefore using Banks you didn't only hold your money in a secure area, but you also have earned more, just by making a saving account.

OpenStudy (nikki_skellington):

omg your a lifesaver thx so much!

OpenStudy (will.h):

yw :)